OpenStudy (anonymous):

need help on proving that the limit of x^4cos 2/x = 0 as x approaches 0

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}x^4\cos\frac{ 2 }{ x }=0\]

OpenStudy (cwrw238):

cos x is bound between -1 and 1 limit as x approaches 0 of x^4 is 0 thus limit of your function as x --> is 0